Write an equation of containing the given point and the parallel to the given line. (3,-8); 9x-5y=4 The equation of the line is y=? I am needing to know how to set this equation up.
Answer by stanbon(75887)

Write an equation of containing the given point and the parallel to the given line. (3,-8); 9x-5y=4 The equation of the line is y=? I am needing to know how to set this equation up.
-----------------
Write the given line in slope-intercept form:
y = (9/5)x - (4/5)
slope is 9/5
----------------
Equation you want has slope 9/5 and y=-8 when x = 3
-8 = (9/5)*3 + b
-8 = 27/5 + b
b = -67/5
=============
EQUATION:
y = (9/5)x - (67/5)